OpenStudy (anonymous):

solve and check 5+13t-4= 6t + -46 - 2t

OpenStudy (anonymous):

5 + 13t - 4 = 6t + -46 - 2t 13t -6t + 2t = -46 - 5 + 4 9t = -47 t = -47 / 9 check yourself....